OpenStudy (anonymous):

What are the cube roots of -1+i sqrt 3

OpenStudy (anonymous):

\[z^3=-1+i\sqrt3=(r(\cos\theta+i\sin\theta))^3\] You have that \[r=\sqrt{(-1)^2+(\sqrt3)^2}=2\] and \[ \begin{cases}2\cos\theta=-1\\ 2\sin\theta=\sqrt3\end{cases}~~\Rightarrow~~\theta=\frac{2\pi}{3}\] The roots will have the form \[z=(-1+\sqrt3)^{1/3}=2^{1/3}\left(\cos\frac{\theta+2k\pi}{3}+i\sin\frac{\theta+2k\pi}{3}\right)\] for \(k=0,1,2\).